Thisis my final problem of the night. hope this problem is beneficial as the others problems given,
This is it: Two truckers leave from the same point driving in opposite directions the first trucker driving at 59 mph and the second at 56mph. The first trucker has a one hour head start. How long will they be able to talk on their car phone if the phone has a 290 mile range.

Answer by mananth(16949)

Trucker I 56 mph 1 hour early
Trucker II 59 mph

Distance 290 miles
Trucker I has traveled 56 miles before Trucker II starts
When Trucker II starts distance between them = (290-56)=234 miles
Let time they take to be be away 234 miles be x hours

Time taken to meet is same
t=d/r
combined speed= (56+59)= 115 mph

115 x = 234
/115
x= 2.03 hours
They can talk for 2.03 hours